OpenStudy (anonymous):

Ralph is 3 times as old as Sara. In 6 years, Ralph will be only twice as old as Sara will be then. Find Ralph's age now. 6 12 18 24

OpenStudy (anonymous):

let Sara's age = x. Then Ralph's age now = 3x. 3x+6= 2(x+6) 3x+6 =2x+12 x =6 (Sara's age now) Ralph is 3x=18 now